# Robotic Learning for Informative Path Planning

### v1

@@ 4.3 Imitation Learning
A simple form of imitation learning (IL) can be thought of as a supervised learning problem from Sec. 4.1 typically referred to as behavioral cloning Pomerleau (1991). The objective is to train a stochastic policy $\pi:\mathcal{S}\to\mathcal{A}$ where the policy $\pi_{\mathbf{\theta}}$ is parameterized by $\mathbf{\theta}$. The training objective is to maximize the likelihood of actions from a dataset $\mathcal{D}$ of expert state-action pairs:
@@
IL has been effectively applied in robotic scenarios, including indoor autonomous exploration, path planning, and multi-robot coordination. Liu et al. (2023) introduce the Learning to Explore (L2E) framework, combining IL with deep RL for UGVs in indoor environments. This approach pretrains exploration policies using IL and then refines them with RL, leading to improved sample efficiency and training speed and outperforming heuristic methods in diverse environments.
Choudhury et al. (2018) propose a data-driven IL framework to train planning policies by imitating a clairvoyant oracle with ground truth knowledge of the world map. This approach efficiently trains policies based on partial information for tasks like AIPP and motion planning, showing significant performance gains over state-of-the-art algorithms, even in real UAV applications. Reinhart et al. (2020) develop a learning-based path planning approach for UAVs in unknown subterranean environments. They use a graph-based path planner as a training expert for IL, resulting in a policy that guides autonomous exploration with reduced computational costs and less reliance on consistent, online reconstructed maps.
Li et al. (2020) address decentralized multi-robot path planning by proposing a combined model that synthesizes local communication and decision-making policies. Their architecture, comprising a convolutional neural network and a graph neural network, was trained to imitate an expert. This model demonstrates its effectiveness in decentralized planning with local communication and observations in cluttered workspaces. Tzes et al. (2023) tackles multi-robot AIPP with their Information-aware Graph Block Network (I-GBNet). Trained via IL, this network aggregates information over a communication graph and offers sequential decision-making in a distributed manner. Its scalability and robustness were validated in large-scale experiments involving dynamic target tracking and localization. Lastly, Dai et al. (2022) explores indoor environments using an RGB-D camera and generative adversarial imitation learning (GAIL). They modified the ORB-SLAM2 method for enhanced navigation and trained a camera view planning policy with GAIL, yielding efficient exploration and reducing tracking failure in indoor environments.
@@ 4.4 Active Learning
The problem of AIPP can also be seen as an active learning (AL) process, whereby the learner uses information measures to choose actions for collecting useful or descriptive data (Taylor et al., 2021). Here, we mathematically connect the general concept of AL to robotic AIPP theory. We then draw parallels between different components of AL and the learning-based AIPP methods presented in the previous subsections. Lastly, we elaborate on an emerging line of work in AIPP that uses AL objectives to improve sensor models in new environments.
AL addresses the task of finding the most informative data in a set of unlabeled samples such that a model is maximally improved if that sample is labeled and added to the existing training dataset (Settles, 2009). We define the AL problem as follows:
@@
In the classical environmental monitoring or exploration problem setup, the environment $\xi$ is characterized by the set of available queries $\mathcal{Q}$, e.g., robot poses $q_{i}=\mathbf{x}_{i}\in\xi$. The feature space $\mathcal{F}$ encompasses the set of targets $\mathcal{L}$ establishing a correspondence between feature mappings $F:\xi\to\mathcal{F}$ and the model $\theta$. Further, we cannot directly quantify the true effect of a sequence of queries $q$ on the model performance $P(\theta)$ as this requires re-training on not-yet-acquired oracle-generated targets. Instead, similar to the belief update of $\tilde{F}$ over the true feature mapping $F$ in Eq. 2, at each iteration $i$, AL methods leverage already collected queries $\{q_{1},\ldots,q_{i}\}$ and targets $\{O(q_{1}),\ldots,O(q_{i})\}$ to update the model $\theta$ and ensure maximally informed sequential query selection. We establish this theoretical connection between AIPP and the AL problem to show that the AIPP methods introduced in Sec. 4.1, 4.2 and 4.3 can be viewed as methods for AL in the context of embodied robotic applications. In the following, we identify the models, queries, oracles, and acquisition functions used in different learning-based approaches from an AL perspective.
The model could have any parametric or non-parametric form. As discussed in Sec. 4.1, common models for learning a spatial representation $F:\xi\to\mathcal{F}$ of a feature space $\mathcal{F}$ in an environment $\xi$ are non-parametric Gaussian processes or implicit neural representations to probabilistically model $O:\xi\to\mathcal{F}$. In this case, a query $q_{i}$ is defined by a specific location or area in the environment $q_{i}=\mathbf{x}_{i}\in\xi$, and the oracle’s answer $O(\mathbf{x}_{i})=z_{i}$ is typically given by collecting a new potentially noisy measurement $z_{i}\in\mathcal{F}$ at the location $\mathbf{x}_{i}$ based on the sensor readings. As summarized in Table 2, the performance metric $P$, i.e., evaluation metric, and thus acquisition function $A$, i.e., information criterion $I$, varies with the task. Reductions in the covariance matrix trace of a Gaussian process model or model uncertainty of an implicit neural representation are popular acquisition functions used in exploration and monitoring tasks. Other supervised approaches learn the acquisition function $A$ with deep neural networks (Hepp et al., 2018; Bai et al., 2017; Ly and Tsai, 2019) or directly predict the next best query $q_{i+1}^{*}$, e.g., a view pose to reconstruct a 3D object (Vasquez-Gomez et al., 2021; Mendoza et al., 2020).
@@
To address this problem, a line of recent research considers using AIPP to actively gather data to improve a sensor model for robotic perception (Rückin et al., 2023, 2022b; Blum et al., 2019; Zurbrügg et al., 2022; Chaplot et al., 2021). In this problem setup, the information criterion $I$ in Eq. 1 is defined in terms of how potential future raw sensor readings, e.g., RGB images at certain environment locations, and their associated oracle-generated targets, e.g., semantic segmentation labels, impact the sensor performance. To solve this problem, the key idea is to link the AIPP objective to ideas from AL.
@@
To date, this problem has only been studied in the context of pixel-wise semantic segmentation from RGB images (Rückin et al., 2023, 2022b; Blum et al., 2019; Zurbrügg et al., 2022; Chaplot et al., 2021). These approaches can be divided into two categories depending on how they implement the oracle for labeling new data: self-supervised methods exploit pseudo labels rendered from an online-built semantic map as an oracle, without considering a human annotator, while fully supervised methods only exploit dense pixel-wise human annotations as an oracle. In the first category, self-supervised methods can be used to fuse 2D semantic predictions from different viewpoints into a semantic 3D map (Zurbrügg et al., 2022; Chaplot et al., 2021). After a mapping mission is completed, the semantic 3D map is used to automatically render consistent 2D pseudo labels for re-training. In this line of work, Zurbrügg et al. (2022) propose an AIPP approach to plan viewpoints with high training data novelty for improved pseudo label quality. Similarly, Chaplot et al. (2021) train an RL agent to target low-confidence parts of a 3D map. Although these approaches do not require a human oracle Zurbrügg et al. (2022); Chaplot et al. (2021), they still rely on large labeled pre-training datasets to generate high-quality pseudo labels in new scenes. In contrast, fully supervised methods are more applicable to varying domains and environments. Blum et al. (2019) locally plan paths of high training data novelty. Rückin et al. (2022b) introduce a map-based planning framework supporting multiple acquisition functions (Rückin et al., 2023). However, these methods (Blum et al., 2019; Rückin et al., 2022b, 2023) require considerable human labeling effort to improve robotic vision.
@@
These works highlight the versatility of applying AL methods to robotic learning through AIPP. The combination of AL and AIPP methods demonstrates a principled approach to integrating learning-based components and classically used uncertainty quantification techniques for AL. This integration represents a promising pathway towards robot deployment in more diverse environments with minimal human supervision. However, to harness the potential of AL and AIPP, future research must address multiple open problems. Existing uncertainty quantification methods for deep learning models are known to be overconfident (Abdar et al., 2021) and not well-calibrated in out-of-distribution scenarios. Further, fully and self-supervised approaches for sensor model improvement are limited in terms of human annotation effort and applicability to out-of-distribution environments, respectively. One possible direction is to take the strengths of both approaches and develop hybrid semi-supervised approaches (Rückin et al., 2024). We may be able to deploy more versatile robots by combining AIPP methods for environment monitoring and exploration (Sec. 4.1 to Sec. 4.3) with methods for targeted training data collection (Rückin et al., 2023, 2022b; Blum et al., 2019; Zurbrügg et al., 2022; Chaplot et al., 2021).
@@ 5 Applications
Table 5 overviews AIPP applications in robotics and describes their relevant practical aspects. Our taxonomy considers four broad application areas: (i) environmental monitoring; (ii) exploration and search; (iii) semantic scene understanding; and (iv) active SLAM. For reference, in the last column, we also include the learning method: supervised learning (‘SL’); reinforcement learning (‘RL’); imitation learning (‘IL’); and/or active learning (‘AL’).
@@
In general, our survey illustrates that the works are broadly applied in terrestrial, aquatic, and aerial domains and on diverse robot platforms. AIPP methods organically find usage in environmental monitoring scenarios to gather data about physical parameters, e.g., signal strength (Hollinger and Sukhatme, 2014; Denniston et al., 2023a; Bayerlein et al., 2021; Denniston et al., 2023b; Yang et al., 2023a; Song et al., 2015), temperature (Meliou et al., 2007; Rückin et al., 2022a; Westheider et al., 2023), or elevation (Viseras and Garcia, 2019; Zacchini et al., 2023). In particular, AIPP enables efficient data collection in large environments using sensors with limited coverage area, such as point-based sensors. For mapping, Gaussian processes are most commonly used to capture spatial correlations found in natural phenomena that are measured by a continuous variable. Several recent works use RL to learn the best data-gathering actions (Rückin et al., 2022a; Cao et al., 2023a; Denniston et al., 2023b), including achieving coordinated monitoring with multi-robot teams (Westheider et al., 2023; Zhu et al., 2023; Yanes Luis et al., 2023; Bayerlein et al., 2021).
@@
AIPP strategies for exploration and search can be further broken down in terms of two distinct tasks: exploration of bounded unknown environments, i.e., outwards-facing scenarios, and active reconstruction of unknown objects or scenes, i.e., inwards-facing scenarios. The former necessarily accounts for possible obstacles in unknown regions and typically relies on depth-based sensing to construct volumetric maps as an environment is explored (Zwecher et al., 2022; Lodel et al., 2022; Choudhury et al., 2018; Tzes et al., 2023; Cao et al., 2023b). In contrast, most approaches for object reconstruction (Vasquez-Gomez et al., 2021; Lee et al., 2022; Jin et al., 2023; Sünderhauf et al., 2023; Mendoza et al., 2020) assume obstacle-free environments with a constrained hemispherical action space for a robot camera around the target object of interest. While NeRF-based methods are gaining popularity for this task (Jin et al., 2023; Sünderhauf et al., 2023; Lee et al., 2022; Yan et al., 2023; Zhang et al., 2023; Pan et al., 2022, 2023b), an open research challenge is applying them for exploration in more complex environments involving multiple objects and clutter.
@@
Only a few works integrate semantic information or consider localization uncertainty in planning objectives for learning-based AIPP. Several approaches (Blum et al., 2019; Georgakis et al., 2022b; Zurbrügg et al., 2022; Rückin et al., 2023, 2022b, 2024) leverage semantics to improve deep learning-based sensor models via AL. Although semantic scene understanding provides valuable cues for learning complex AIPP behavior (Schmid et al., 2023), a key development barrier is the lack of high-quality semantic datasets relevant to the domains of robotic AIPP applications. The common assumption of perfect localization uncertainty represents a significant limitation to the applicability of existing methods, particularly in outdoor environments which may rely on imprecise Global Positioning System (GPS) signals (Caley et al., 2019; Hitz et al., 2017; Westheider et al., 2023; Choudhury et al., 2020; Hüttenrauch et al., 2019). Many approaches also neglect sensor uncertainty and noise, especially in environmental monitoring scenarios where models for specialized sensors are scarce. This is a major drawback since imperfect sensing has been shown to have significant effects on the quality of collected data and resulting planning strategy (Popović et al., 2020; Arora et al., 2019; Zacchini et al., 2023).
@@
Our taxonomy reveals a gap in methods accounting for robot motion constraints while planning, with some methods assuming that the robot sensor can teleport between locations (Jin et al., 2023; Pan et al., 2022; Sünderhauf et al., 2023) or perform only incremental motions relative to its current position (Westheider et al., 2023; Viseras et al., 2021; Bayerlein et al., 2021). One open issue is the lack of guarantees on dynamic feasibility, which is essential for long-term path planning with agile, fast platforms such as small UAVs. Although there has been research on multi-robot missions (Westheider et al., 2023; Zhu et al., 2023; Yang et al., 2023a; Gao and Zhang, 2022; Hüttenrauch et al., 2019), there are several areas that have not been widely studied, such as tasks involving multi-sensor fusion (Ott et al., 2023), temporal changes in the environment (Caley et al., 2019; Marchant et al., 2014), and combining multiple problems simultaneously, e.g., semantic scene understanding and active SLAM.
@@
Environmental monitoring
(Velasco et al., 2020)
Air quality monitoring
UAV
Point-based gas sensor
2D
Mapping
(Choi and Cielniak, 2021)
Air quality monitoring
UGV
Point-based sensor
2D
Mapping
SL, RL
(Viseras and Garcia, 2019)
Elevation mapping
UAV team
Ultrasound height sensor
2D
Mapping
SL, RL
(Zeng et al., 2022)
Fruit monitoring
Robot arm
RGB-D camera
3D
Mapping
ROS $*{}^{*}$
RL
(Menon et al., 2023)
Fruit monitoring
Robot arm
RGB-D camera
3D
Mapping
ROS $*{}^{*}$
SL
(Denniston et al., 2023b)
Intensity monitoring
AUV/UGV/UAV
@@
Generic camera, point-based sensor
2D/3D
Mapping
SL, RL
(Cao et al., 2023a)
Intensity monitoring
UGV
Point-based sensor
2D
Mapping
SL, RL
(Yang et al., 2023a)
Intensity monitoring
UGV team
Point-based sensor
2D
Mapping
SL, RL
(Song et al., 2015)
Intensity monitoring
UAV
Point-based sensor
2D
Mapping, sensing
Nonholonomic constraints
SL
(Hitz et al., 2017)
Lake bacteria monitoring
AUV
@@
Point sensor for phycoerythrin fluorescence
3D
Mapping
Spline path
SL
(Duecker et al., 2021)
Ocean pollution monitoring
AUV
@@
Pressure sensor, general water characterization
2D
Mapping
Spline path
RL
(Popović et al., 2020)
Precision agriculture
UAV
Generic camera
3D
Mapping, sensing, semantics
Spline path
SL
(Rückin et al., 2022a)
Precision agriculture
UAV
Generic camera
3D
Mapping, sensing
SL, RL
(Westheider et al., 2023)
Precision agriculture
UAV team
Generic camera
3D
Mapping, sensing
RL
(Arora et al., 2019)
Scientific data gathering
UGV
@@
Generic camera, ultraviolet light source, laser scanner
2D
Mapping, sensing
Omnidirectional drive
RL
(Hollinger and Sukhatme, 2014)
Signal strength monitoring
AUV
Point-based sensor
2D
Mapping
SL
(Denniston et al., 2023a)
Signal strength monitoring
AUV/UGV
@@
Generic camera, point-based sensor
2D
Mapping, path executability
RL
(Wei and Zheng, 2020)
Signal strength monitoring
UGV
Point-based sensor
2D
Mapping
SL, RL
(Caley et al., 2019)
Spatiotemporal field monitoring
AUV/UGV
Point-based sensor
2D
Mapping
SL
(Marchant et al., 2014)
Spatiotemporal field monitoring
UGV
Point-based sensor
2D
Mapping
Spline path
SL, RL
(Tzes et al., 2023)
Target tracking and localization
UGV team
RGB-D sensor
2D
Mapping, sensing
IL
(Meliou et al., 2007)
Temperature monitoring
Sensor network
Generic
2D
Mapping
SL
(Vivaldini et al., 2019)
Tree disease classification
UAV
RGB camera
2D
Mapping, semantics
SL
(Zacchini et al., 2023)
Underwater mapping
AUV
Sonar
3D
Mapping, sensing
Dubins vehicle model
SL
(Yanes Luis et al., 2023)
Water quality monitoring
AUV team
Water quality sensor
2D
Mapping
SL, RL
(Viseras et al., 2021)
Wildfire monitoring
UAV team
Thermal camera
2D
RL
(Bayerlein et al., 2021)
Wireless data harvesting
UAV team
Signal receptor
2D
Signal emissions
RL
Exploration and search
(Dai et al., 2022)
Camera view planning
UGV
RGB-D camera
3D
IL
(Yang et al., 2023b)
Exploration and landmark localization
UAV
Depth sensor
2D/3D
Mapping
RL
(Zhu et al., 2023)
Exploration and target search
UGV team
Depth sensor
2D
RL
(Choudhury et al., 2018)
Exploration, search-based planning
UGV/UAV
RGB-D camera
2D
Mapping
IL
(Lodel et al., 2022)
Floor plan exploration
UAV
Laser
2D
Mapping, sensing
Unicycle model
RL
(Liu et al., 2023)
Indoor exploration
UGV
Depth sensor
2D
RL, IL
(Zwecher et al., 2022)
Indoor exploration
UGV
Depth sensor
2D
Map prediction
SL, RL
(Bai et al., 2017)
Indoor exploration
UGV
Depth sensor
2D
Mapping, sensing
SL
(Tao et al., 2023)
Indoor exploration
UAV
Depth sensor
3D
Mapping, sensing
Spline path
SL
(Cao et al., 2023b)
Indoor exploration
UGV
Depth sensor
2D
Mapping, sensing
Local $*{}^{*}$ (Cao et al., 2022)
RL
@@
    *** Motion constraints incorporated only during plan execution.
@@
Exploration and search
(Niroui et al., 2019)
Indoor exploration
UGV
Depth sensor
2D
Mapping, sensing
ROS $*{}^{*}$
RL
(Chen et al., 2019)
Indoor exploration
UGV
Depth sensor
2D/3D
Mapping, sensing
RL
(Shrestha et al., 2019)
Indoor exploration
UGV
Laser scanner
2D
Mapping
SL
(Li et al., 2023)
Indoor exploration
UGV
RGB-D camera
2D
@@
Local planner using Fast Marching Method
SL
(Schmid et al., 2022)
Indoor exploration
UGV
RGB-D camera
2D
Mapping, sensing
SL
(Schmid et al., 2023)
Indoor exploration
UAV
RGB-D camera
3D
@@
Mapping, sensing, scene completion
Velocity ramp model
SL
(Georgakis et al., 2022a)

### v3

@@ 4.3 Imitation Learning
A simple form of imitation learning (IL) can be thought of as a supervised learning problem from Sec. 4.1 typically referred to as behavioral cloning [131]. The objective is to train a stochastic policy $\pi:\mathcal{S}\to\mathcal{A}$ where the policy $\pi_{\mathbf{\theta}}$ is parameterized by $\mathbf{\theta}$. The training objective is to maximize the likelihood of actions from a dataset $\mathcal{D}$ of expert state-action pairs:
@@
IL has been effectively applied in robotic scenarios, including indoor autonomous exploration, path planning, and multi-robot coordination. Liu et al. [46] introduce the Learning to Explore (L2E) framework, combining IL with deep RL for UGVs in indoor environments. This approach pretrains exploration policies using IL and then refines them with RL, leading to improved sample efficiency and training speed and outperforming heuristic methods in diverse environments.
Choudhury et al. [31] propose a data-driven IL framework to train planning policies by imitating a clairvoyant oracle with ground truth knowledge of the world map. This approach efficiently trains policies based on partial information for tasks like AIPP and motion planning, showing significant performance gains over state-of-the-art algorithms, even in real UAV applications. Reinhart et al. [42] develop a learning-based path planning approach for UAVs in unknown subterranean environments. They use a graph-based path planner as a training expert for IL, resulting in a policy that guides autonomous exploration with reduced computational costs and less reliance on consistent, online reconstructed maps.
Li et al. [106] address decentralized multi-robot path planning by proposing a combined model that synthesizes local communication and decision-making policies. Their architecture, comprising a convolutional neural network and a graph neural network, was trained to imitate an expert. This model demonstrates its effectiveness in decentralized planning with local communication and observations in cluttered workspaces. Tzes et al. [74] tackles multi-robot AIPP with their Information-aware Graph Block Network (I-GBNet). Trained via IL, this network aggregates information over a communication graph and offers sequential decision-making in a distributed manner. Its scalability and robustness were validated in large-scale experiments involving dynamic target tracking and localization. Lastly, Dai et al. [55] explores indoor environments using an RGB-D camera and generative adversarial imitation learning (GAIL). They modified the ORB-SLAM2 method for enhanced navigation and trained a camera view planning policy with GAIL, yielding efficient exploration and reducing tracking failure in indoor environments.
@@ 4.4 Active Learning
The problem of AIPP can also be seen as an active learning (AL) process, whereby the learner uses information measures to choose actions for collecting useful or descriptive data [15]. Here, we mathematically connect the general concept of AL to robotic AIPP theory. We then draw parallels between different components of AL and the learning-based AIPP methods presented in the previous subsections. Lastly, we elaborate on an emerging line of work in AIPP that uses AL objectives to improve sensor models in new environments.
AL addresses the task of finding the most informative data in a set of unlabeled samples such that a model is maximally improved if that sample is labeled and added to the existing training dataset [132]. We define the AL problem as follows:
@@
In the classical environmental monitoring or exploration problem setup, the environment $\xi$ is characterized by the set of available queries $\mathcal{Q}$, e.g., robot poses $q_{i}=\mathbf{x}_{i}\in\xi$. The feature space $\mathcal{F}$ encompasses the set of targets $\mathcal{L}$ establishing a correspondence between feature mappings $F:\xi\times\mathbb{R}\to\mathcal{F}$ and the model $\theta$. Further, we cannot directly quantify the true effect of a sequence of queries $q$ on the model performance $P(\theta)$ as this requires re-training on not-yet-acquired oracle-generated targets. Instead, similar to the belief update of $\tilde{F}$ over the true feature mapping $F$ in Eq. 2, at each iteration $i$, AL methods leverage already collected queries $\{q_{1},\ldots,q_{i}\}$ and targets $\{O(q_{1}),\ldots,O(q_{i})\}$ to update the model $\theta$ and ensure maximally informed sequential query selection. We establish this theoretical connection between AIPP and the AL problem to show that the AIPP methods introduced in Sec. 4.1, 4.2 and 4.3 can be viewed as methods for AL in the context of embodied robotic applications. In the following, we identify the models, queries, oracles, and acquisition functions used in different learning-based approaches from an AL perspective.
The model could have any parametric or non-parametric form. As discussed in Sec. 4.1, common models for learning a spatio(temporal) representation $F$ of a feature space $\mathcal{F}$ in an environment $\xi$ are non-parametric Gaussian processes or implicit neural representations to probabilistically model $O:\xi\to\mathcal{F}$. In this case, a query $q_{i}$ is defined by a specific location or area in the environment $q_{i}=\mathbf{x}_{i}\in\xi$, and the oracle’s answer $O(\mathbf{x}_{i})=z_{i}$ is typically given by collecting a new potentially noisy measurement $z_{i}\in\mathcal{F}$ at the location $\mathbf{x}_{i}$ based on the sensor readings. As summarized in Table 2, the performance metric $P$, i.e., evaluation metric, and thus acquisition function $A$, i.e., information criterion $I$, varies with the task. Reductions in the covariance matrix trace of a Gaussian process model or model uncertainty of an implicit neural representation are popular acquisition functions used in exploration and monitoring tasks. Approaches leveraging a well-defined probabilistic map belief as a model and maximising an information-theoretically motivated acquisition function are often closely related to the theory of optimal experimental design [133]. Examples are AIPP approaches that sequentially update Gaussian process map beliefs aiming to maximize its covariance trace reduction [71, 65, 3, 62, 2] as they are derived from A-optimal experiment design strategies. Other supervised approaches learn the acquisition function $A$ with deep neural networks [56, 28, 107] or directly predict the next best query $q_{i+1}^{*}$, e.g., a view pose to reconstruct a 3D object [47, 48].
@@
To address this problem, a line of recent research considers using AIPP to actively gather data to improve a sensor model for robotic perception [90, 134, 88, 89, 92]. In this problem setup, the information criterion $I$ in Eq. 1 is defined in terms of how potential future raw sensor readings, e.g., RGB images at certain environment locations, and their associated oracle-generated targets, e.g., semantic segmentation labels, impact the sensor performance. To solve this problem, the key idea is to link the AIPP objective to ideas from AL.
@@
To date, this problem has only been studied in the context of pixel-wise semantic segmentation from RGB images [90, 134, 88, 89, 92]. These approaches can be divided into two categories depending on how they implement the oracle for labeling new data: self-supervised methods exploit pseudo labels rendered from an online-built semantic map as an oracle, without considering a human annotator, while fully supervised methods only exploit dense pixel-wise human annotations as an oracle. In the first category, self-supervised methods can be used to fuse 2D semantic predictions from different viewpoints into a semantic 3D map [89, 92]. After a mapping mission is completed, the semantic 3D map is used to automatically render consistent 2D pseudo labels for re-training. In this line of work, Zurbrügg et al. [89] propose an AIPP approach to plan viewpoints with high training data novelty for improved pseudo label quality. Similarly, Chaplot et al. [92] train an RL agent to target low-confidence parts of a 3D map. Although these approaches do not require a human oracle [89, 92], they still rely on large labeled pre-training datasets to generate high-quality pseudo labels in new scenes. In contrast, fully supervised methods are more applicable to varying domains and environments. Blum et al. [88] locally plan paths of high training data novelty. Rückin et al. [134] introduce a map-based planning framework supporting multiple acquisition functions [90]. However, these methods [88, 134, 90] require considerable human labeling effort to improve robotic vision.
@@
These works highlight the versatility of applying AL methods to robotic learning through AIPP. The combination of AL and AIPP methods demonstrates a principled approach to integrating learning-based components and classically used uncertainty quantification techniques for AL. This integration represents a promising pathway towards robot deployment in more diverse environments with minimal human supervision. However, to harness the potential of AL and AIPP, future research must address multiple open problems. Existing uncertainty quantification methods for deep learning models are known to be overconfident [135] and not well-calibrated in out-of-distribution scenarios. Further, fully and self-supervised approaches for sensor model improvement are limited in terms of human annotation effort and applicability to out-of-distribution environments, respectively. One possible direction is to take the strengths of both approaches and develop hybrid semi-supervised approaches [91]. We may be able to deploy more versatile robots by combining AIPP methods for environment monitoring and exploration (Sec. 4.1 to Sec. 4.3) with methods for targeted training data collection [90, 134, 88, 89, 92].
Figure 3: Summary statistics for the papers reviewed in Sec. 5.
@@ 5 Applications
Table 5 overviews AIPP applications in robotics and describes their relevant practical aspects. Our taxonomy considers four broad application areas: (i) environmental monitoring; (ii) exploration and search; (iii) semantic scene understanding; and (iv) active SLAM. For reference, in the last column, we also include the learning method: supervised learning (‘SL’); reinforcement learning (‘RL’); imitation learning (‘IL’); and/or active learning (‘AL’). In addition, in Fig. 3, we provide visual summary statistics for our survey according to the application area, learning method, and planning space considered by each paper.
@@
In general, our survey illustrates that the works are broadly applied in terrestrial, aquatic, and aerial domains and on diverse robot platforms. AIPP methods organically find usage in environmental monitoring scenarios to gather data about physical parameters, e.g., signal strength [68, 4, 120, 59, 63, 60], temperature [76, 5, 27], or elevation [6, 45]. In particular, AIPP enables efficient data collection in large environments using sensors with limited coverage area, such as point-based sensors. For mapping, Gaussian processes are most commonly used to capture spatial correlations found in natural phenomena that are measured by a continuous variable. Several recent works use RL to learn the best data-gathering actions [5, 2, 59], including achieving coordinated monitoring with multi-robot teams [27, 136, 64, 120].
@@
AIPP strategies for exploration and search can be further broken down in terms of two distinct tasks: exploration of bounded unknown environments, i.e., outwards-facing scenarios, and active reconstruction of unknown objects or scenes, i.e., inwards-facing scenarios. The former necessarily accounts for possible obstacles in unknown regions and typically relies on depth-based sensing to construct volumetric maps as an environment is explored [54, 32, 31, 74, 41]. In contrast, most approaches for object reconstruction [47, 84, 77, 81, 48] assume obstacle-free environments with a constrained hemispherical action space for a robot camera around the target object of interest. While NeRF-based methods are gaining popularity for this task [77, 81, 84, 83, 97, 78, 82], an open research challenge is applying them for exploration in more complex environments involving multiple objects and clutter.
@@
Only a few works integrate semantic information or consider localization uncertainty in planning objectives for learning-based AIPP. Several approaches [88, 94, 89, 90, 134, 91] leverage semantics to improve deep learning-based sensor models via AL. Although semantic scene understanding provides valuable cues for learning complex AIPP behavior [38], a key development barrier is the lack of high-quality semantic datasets relevant to the domains of robotic AIPP applications. The common assumption of perfect localization uncertainty represents a significant limitation to the applicability of existing methods, particularly in outdoor environments which may rely on imprecise Global Positioning System (GPS) signals [49, 65, 27, 7, 72]. Many approaches also neglect sensor uncertainty and noise, especially in environmental monitoring scenarios where models for specialized sensors are scarce. This is a major drawback since imperfect sensing has been shown to have significant effects on the quality of collected data and resulting planning strategy [3, 126, 45].
@@
Our taxonomy reveals a gap in methods accounting for robot motion constraints while planning, with some methods assuming that the robot sensor can teleport between locations [77, 78, 81] or perform only incremental motions relative to its current position [27, 35, 120]. One open issue is the lack of guarantees on dynamic feasibility, which is essential for long-term path planning with agile, fast platforms such as small UAVs. Although there has been research on multi-robot missions [27, 136, 63, 53, 72], there are several areas that have not been widely studied, such as tasks involving multi-sensor fusion [62], temporal changes in the environment [49, 71], and combining multiple problems simultaneously, e.g., semantic scene understanding and active SLAM.
@@
Environmental monitoring [95] Air quality monitoring UAV Point-based gas sensor 2D Mapping - -
[69] Air quality monitoring UGV Point-based sensor 2D Mapping - SL, RL
[6] Elevation mapping UAV team Ultrasound height sensor 2D Mapping - SL, RL
[43] Fruit monitoring Robot arm RGB-D camera 3D Mapping ROS MoveIt RL
[137] Fruit monitoring Robot arm RGB-D camera 3D Mapping ROS MoveIt SL
@@
[59] Intensity monitoring AUV/UGV/UAV Generic camera, point-based sensor 2D/3D Mapping - SL, RL
[2] Intensity monitoring UGV Point-based sensor 2D Mapping - SL, RL
[63] Intensity monitoring UGV team Point-based sensor 2D Mapping - SL, RL
[60] Intensity monitoring UAV Point-based sensor 2D Mapping, sensing Nonholonomic constraints SL
@@
[65] Lake bacteria monitoring AUV Point sensor for phycoerythrin fluorescence 3D Mapping Spline path SL
@@
[73] Ocean pollution monitoring AUV Pressure sensor, general water characterization 2D Mapping Spline path RL
[3] Precision agriculture UAV Generic camera 3D Mapping, sensing, semantics Spline path SL
[5] Precision agriculture UAV Generic camera 3D Mapping, sensing - SL, RL
[27] Precision agriculture UAV team Generic camera 3D Mapping, sensing - RL
@@
[126] Scientific data gathering UGV Generic camera, ultraviolet light source, laser scanner 2D Mapping, sensing Omnidirectional drive RL
[68] Signal strength monitoring AUV Point-based sensor 2D Mapping - SL
@@
[4] Signal strength monitoring AUV/UGV Generic camera, point-based sensor 2D Mapping, path executability - RL
[66] Signal strength monitoring UGV Point-based sensor 2D Mapping - SL, RL
[49] Spatiotemporal field monitoring AUV/UGV Point-based sensor 2D Mapping - SL
[71] Spatiotemporal field monitoring UGV Point-based sensor 2D Mapping Spline path SL, RL
[74] Target tracking and localization UGV team RGB-D sensor 2D Mapping, sensing - IL
[76] Temperature monitoring Sensor network Generic 2D Mapping - SL
[67] Tree disease classification UAV RGB camera 2D Mapping, semantics - SL
[45] Underwater mapping AUV Sonar 3D Mapping, sensing Dubins vehicle model SL
[64] Water quality monitoring AUV team Water quality sensor 2D Mapping - SL, RL
[35] Wildfire monitoring UAV team Thermal camera 2D - - RL
[120] Wireless data harvesting UAV team Signal receptor 2D Signal emissions - RL
Exploration and search [55] Camera view planning UGV RGB-D camera 3D - - IL
[121] Exploration and landmark localization UAV Depth sensor 2D/3D Mapping - RL
[136] Exploration and target search UGV team Depth sensor 2D - - RL
[31] Exploration, search-based planning UGV/UAV RGB-D camera 2D Mapping - IL
[32] Floor plan exploration UAV Laser 2D Mapping, sensing Unicycle model RL
[46] Indoor exploration UGV Depth sensor 2D - - RL, IL
[54] Indoor exploration UGV Depth sensor 2D Map prediction - SL, RL
[28] Indoor exploration UGV Depth sensor 2D Mapping, sensing - SL
[51] Indoor exploration UAV Depth sensor 3D Mapping, sensing Spline path SL
[41] Indoor exploration UGV Depth sensor 2D Mapping, sensing Local planner [138] RL
@@
    * Motion constraints incorporated only during plan execution.
@@
Exploration and search [40] Indoor exploration UGV Depth sensor 2D Mapping, sensing ROS move_base RL
[33] Indoor exploration UGV Depth sensor 2D/3D Mapping, sensing - RL
[44] Indoor exploration UGV Laser scanner 2D Mapping - SL
@@
[52] Indoor exploration UGV RGB-D camera 2D - Local planner using Fast Marching Method SL
[39] Indoor exploration UGV RGB-D camera 2D Mapping, sensing - SL
@@
[38] Indoor exploration UAV RGB-D camera 3D Mapping, sensing, scene completion Velocity ramp model SL